The width of a rectangle is 2 inches less than its length. Find the length and width if the area of the rectangle is 35 inches wide.
Answer by algebrahouse.com(1659)

x = length
x - 2 = width {width is 2 less than length}

Area = length x width

x(x - 2) = 35 {area is length x width}
x² - 2x = 35 {used distributive property}
x² - 2x - 35 = 0 {subtracted 35 from each side}
(x - 7)(x + 5) = 0 {factored into two binomials}
x - 7 = 0 or x + 5 = 0 {set each factor equal to 0}
x = 7 or x = -5 {solved each equation for x}
x = 7 {length cannot be negative
x - 2 = 5 {substituted 7, in for x, into x - 2}

length is 7 in. and width is 5 in.

For more help from me, visit: www.algebrahouse.com